# Think-While-Generating: On-the-Fly Reasoning for Personalized Long-Form Generation

### v1

@@ 1 Introduction
Large language models (LLMs) have achieved significant advances through preference alignment techniques, enabling them to produce outputs that more closely reflect human expectations (instructGPT; dpo). Yet, common preference alignment (dpo; ppo) primarily focuses on population-level preferences, often overlooking the diverse characteristics and nuanced needs of individual users (PerAlign-survey; PerAlign-survey2; rebuttal_1; rebuttal_2; rebuttal_3; rebuttal_4). These limitations inevitably curb user satisfaction and hinder the more widespread adoption of LLMs in everyday life (PerAlign-survey; iclr-LongMemEval; iclr-PersonalLLM; iclr-PrefEval). To bridge this gap, personalization has emerged as an indispensable component of LLM alignment—an essential step toward integrating broad, generalized utility with the fine-grained, personalized adaptability required for truly user-centric AI (ICML_ACL_keynote). Recently, increasing attention from both academia and industry has been devoted to this direction.
Existing LLM personalization methods typically rely on either customizing user-specific prompts via RAG (RAG1-intro; RAG2-intro; RAG3-intro), or fine-tuning the model (Amulet; PAD; drift) to produce personalized outputs. While straightforward, these approaches often struggle in real-world scenarios where user interests are not explicitly stated but are only implicitly reflected in historical behavior. Such methods lack a mechanism to reason over implicit user preferences, potentially resulting in suboptimal personalization. Aiming at addressing this limitation, a new “think-then-generate” paradigm for LLM personalization has emerged (Reasoning_1; Reasoning_2; Reasoning_3; Reasoning_0): rather than directly generating the final response, the LLM first produces an in-depth analysis of the user’s preferences based on their input, and then generates responses aligned with this analysis. By leveraging the model’s reasoning capabilities, this approach can potentially capture implicit user interests more effectively, enabling more accurate personalization.
@@
To implement the think-while-generating paradigm, a key challenge is efficiency: frequent reasoning across response segments substantially increases both training and inference time as the number of segments grows. Latent reasoning (coconut) offers a more efficient alternative to standard chain-of-thought (CoT) approaches (CoT_1) by compressing the reasoning path into a few hidden-state tokens, avoiding the overhead of generating lengthy CoT sequences. However, directly applying existing latent reasoning methods remains inefficient. Current approaches (reasoning_latent1; reasoning_latent2) typically use the same LLM for reasoning and generation, forcing each output token generation to wait for all preceding reasoning tokens. Non-reasoning training can be accelerated by feeding all ground-truth tokens simultaneously, enabling parallel optimization. With reasoning tokens, this parallelism is lost, as the model must wait for the reasoning outputs, slowing down training. For inference, the time cost also increases with the total number of reasoning tokens. To solve the problem, the key is to keep the parallelism.
@@ 4.1 Experiment Settings
Datasets & Backbone LLMs. We evaluate our FlyThinker on three long-form personalized generation tasks from the user-based split of the LongLamp benchmark (longlamp): Product Review, Abstract Generation, and Topic Writing, which are designed to assess the effectiveness of LLMs in generating personalized, long-text outputs. The model is required to generate a long-form response that aligns with user preferences. Our primary backbone is Qwen2.5-3B-Instruct, with additional experiments on Qwen2.5-7B-Instruct and Gemma-7B-it available in Appendix L and  M.
Baselines & Evaluation Metrics. We compare FlyThinker against the following baselines: (1) Tuning-free methods: query-only generation $(Non-pers$), retrieval-augmented generation $(RAG$), and context steering $(CoS$); (2) Tuning-based methods: supervised fine-tuning $(SFT$), summary-enhanced training (LLM-TRSR(LLM-trsr), causal preference modeling (NextQuill(Nextquill), chain-of-thought reasoning (CoT(CoT_1; CoT_2), and latent reasoning (Coconut(coconut). We report ROUGE-1, ROUGE-L, BLEU, and METEOR scores. Baseline details and results on other evaluation metrics are provided in the Appendix D and E, respectively.
@@ 5 Related Work
Parameter-level personalization. This paradigm configures user-specific parameter modules and fine-tunes them on personal data to achieve individualized effects (related-HYDRA; related-oppu; related-Per-Pcs; related-PROPER). However, due to the inherent sparsity of user data, such models often struggle to achieve robust personalized performance; moreover, the user-isolated training paradigm lacks cross-user generalization and introduces significant maintenance costs in large-scale systems.
Context-level personalization. This line of research focuses on injecting user information into the input context to enable personalization. Two main strategies have emerged. The first, explicit context (PAD; Amulet; drift), where preference information (e.g., “formal style”) is directly provided in the input to guide generation. The second, implicit context (related-implicit-icml; related-implicit-nips), where the LLM infers latent preferences from users’ historical behaviors and interactions.
Reasoning-based personalization. Reasoning-based approaches have gained traction, for instance by incorporating CoT or latent reasoning frameworks to enhance the capture of implicit preferences (related-reasoning-1; related-reasoning-2; reasoning_latent1; reasoning_latent2). Nevertheless, in long-form text generation tasks, personalization is inherently complex and dynamically evolving, making single-pass static reasoning insufficient for capturing users’ shifting needs.
@@ Appendix C Implementation Details
In our study, we fine-tune the Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct model to validate our approach. All experiments, including FlyThinker and baseline models, are conducted on four NVIDIA A100 GPUs. The LLMs are loaded via HuggingFace Transformers (huggingface), and the training is implemented in Python 3.10 using the Verl training framework (verl).
@@ Appendix H Latent Reasoning Trajectory: Visualization and Analysis
Figure 8: Fixing the Task, Varying the User. The sample texts illustrate how reasoning trajectories that are close or distant in the visualization space correspond to differences in generated content.
@@
To examine whether these trajectories relate to the preference, we used GPT-5.1 to compare the samples. We selected the two samples whose users had the most similar trajectory directions (sample0 and sample4) and the one farthest away (sample2) in Figure 7. As shown in Figure 8, GPT-5.1 judged sample4 to be most similar to sample0 in paragraph structure, narrative style, and analytical approach. Sample2 was the least similar. These results show that different users induce distinct reasoning trajectories, and that closer trajectories correspond to more similar writing styles and semantic structures, etc. This suggests that latent reasoning tokens may contain interpretable semantic information useful for user clustering and preference visualization.
@@ Appendix I Reasoner–Generator Relative Size Analysis in FlyThinker
In this section, we added an LLM-based evaluation to provide a more comprehensive assessment of personalized generation quality. Specifically, we use GPT-4o as an evaluator and compare the outputs produced by FlyThinker with those from the strongest baseline, SFT, following a structured evaluation rubric 10. As shown in the Table 5, FlyThinker consistently achieves higher scores across all datasets under the LLM-based evaluation. These findings indicate that FlyThinker’s personalization gains are reflected not only in traditional automatic metrics but also in the LLM judge’s assessments.
@@
Appendix Q Case Studies of Personalized Outputs
In this section, we offer a more intuitive understanding of how FlyThinker enhances personalization quality through several case studies. These examples demonstrate the effect of varying the Reasoner size. For complete samples, please refer to Figure 12.
Question Generate the review…for a product with description “Based on… breaks the world one-hour record on …”. The … is “… more about a battle with depression…”.
User’s Real Writing …… using a home-made bicycle (even using parts from a washing machine) (1) breaks the worldś endurance record in a time trial that had cyclists everywhere in awe.(2) ……, takes Obree to the point of suicide, and a long wrestle with manic depression…….It was a painful lack of self-confidence instilled by years of bullying and by precious little help from his own father, a policeman. (3) As with many trues stories of depression,(4) ……If youŕe looking for an exciting cycle race movie, no, this is not the one for you, but if you want a movie that takes you into the sometimes dark world of the human soul, be prepared for one tough ride.(5) ……As this inspirational story shows: a world record didnt́ satisfy Obreeś demons.(6) ……
Analysis: The user’s real writing includes the following elements: factual information, such as (1) bicycle production and (2) awards; emotional tone, such as (3), which is predominantly serious and somber, in line with the story’s theme; writing techniques, including analogy (4), contrast (5), and (6); and core perspectives, such as (4), which emphasizes that the central focus of the narrative is the protagonist’s internal struggle rather than a competition, and (5), which underscores that the protagonist’s inner world holds more significance than external honors.
FlyThinker (R=3B) …… broke the world one-hour cycling record (1) ……It’s not an Olympic film: it focuses on the battle between Graeme and depression (2)……Graeme is a complex and fascinating human being who has a remarkable ability to focus and achieve, but he has a fragile and troubled mind.(3) …… There are parallels to someone like Lance Armstrong (who also battled depression as a result of an accident.)(4) This movie is not about racing or cycling …… it’s about the pursuit of excellence - but it’s more about the pursuit of mental sanity.(5)
Analysis: The generated content contains many characteristics consistent with the user’s real writing: it incorporates key factual details (1), and maintains an emotional tone that is serious, somber, and centered on depression (3). Additionally, the writing techniques are consistent, utilizing contrast (2), (3), (5), and analogy (3). Furthermore, the core perspectives are aligned, such as (2), which highlights that the story revolves around the protagonist’s internal conflict rather than a competition, and (5), which emphasizes the primacy of the protagonist’s inner world over external honors.
FlyThinker (R=1.5B) Graeme Obree was a Scottish cyclist who, in the mid-1990s, made one of the most remarkable achievements in cycling history.(1) …… He did it using a bicycle that hed́ invented himself (2)…… Butthe movie is about more than this incredible achievement. Itś about Graemeś mental breakdown - when he lost his mind - and his subsequent struggle to come back from the brink.(3) …… But then there are moments of levity too - as we see Graeme with friends or family, cycling around or in his shop, talking to them.(4)……
Analysis: The generated content also includes elements consistent with the user’s writing: it references similar factual information (1) and (2), and employs a consistent writing technique, such as contrast (3). Moreover, the core perspective remains intact, as (3) stresses the internal struggle of the protagonist rather than focusing on a competition. However, while the emotional tone is predominantly serious and somber, it occasionally incorporates unrelated emotional tones, such as (4).
FlyThinker (R=0.5B) …… this story of one manś extraordinary battle with depression.(1)…… a frame of welded together steel pipes, some washers, two broken front wheels, and one broken pedal.(2) …… I had seen some photos of Graeme cycling around Glasgow, and I had read the article about his cycling career in the Sunday Herald back in the late 1980s.(3) …… One small gripe: the ending of the film, which is a replay of the race, feels too abrupt.(4)……
Analysis: The generated content contains some features that align with the user’s writing, such as mentioning similar factual information (2) and introducing the theme of the story, as in (1), which emphasizes the protagonist’s internal struggle rather than a competition. However, the content also includes a significant amount of irrelevant information (3), (4), and does not maintain the serious, somber emotional tone. Instead, it adopts a lighter emotional approach (3) and fails to employ the user’s typical writing techniques.
Our results show that larger Reasoner models better capture the user’s tone, emotional depth, and narrative structure better.
Appendix R Difficulty-Aware Evaluation for Personalized Tasks
@@
Figure 12: Case study examples comparing FlyThinker under different Reasoner scales.

### v3

@@ 1 Introduction
Large language models (LLMs) have achieved significant advances through preference alignment techniques, enabling them to produce outputs that more closely reflect human expectations (Ouyang et al., 2022; Rafailov et al., 2023). Yet, common preference alignment (Rafailov et al., 2023; Schulman et al., 2017) primarily focuses on population-level preferences, often overlooking the diverse characteristics and nuanced needs of individual users (Guan et al., 2025; Xie et al., 2025; Li et al., 2024; Zhao et al., 2025b; Qiu et al., 2025a; b; Zhao et al., 2026; Zhang et al., 2025c). These limitations inevitably curb user satisfaction and hinder the more widespread adoption of LLMs in everyday life (Guan et al., 2025; Wu et al., 2025; Zollo et al., 2025; Zhao et al., 2025a; Wang et al., 2025a; Shi et al., 2024; Wang et al., 2025b). To bridge this gap, personalization has emerged as an indispensable component of LLM alignment—an essential step toward integrating broad, generalized utility with the fine-grained, personalized adaptability required for truly user-centric AI (Sorensen et al., 2024).
Existing LLM personalization methods typically rely on either customizing user-specific prompts via RAG (Qian et al., 2024; Mysore et al., 2023b; Zhang et al., 2025a), or fine-tuning the model (Zhang et al., 2025d; Chen et al., 2025; Kim et al., 2025c) to produce personalized outputs. While straightforward, these approaches often struggle in real-world scenarios where user interests are not explicitly stated but are only implicitly reflected in historical behavior. Such methods lack a mechanism to reason over implicit user preferences, potentially resulting in suboptimal personalization. Aiming at addressing this limitation, a new “think-then-generate” paradigm for LLM personalization has emerged (Luo et al., 2025b; Kim et al., 2025a; Li et al., 2025b; Salemi et al., 2025c): rather than directly generating the final response, the LLM first produces an in-depth analysis of the user’s preferences based on their input, and then generates responses aligned with this analysis. By leveraging the model’s reasoning capabilities, this approach can potentially capture implicit user interests more effectively, enabling more accurate personalization.
@@
To implement the think-while-generating paradigm, a key challenge is efficiency: frequent reasoning across response segments substantially increases both training and inference time as the number of segments grows. Latent reasoning (Hao et al., 2024) offers a more efficient alternative to standard chain-of-thought (CoT) approaches (Wei et al., 2022) by compressing the reasoning path into a few hidden-state tokens, avoiding the overhead of generating lengthy CoT sequences. However, directly applying existing latent reasoning methods remains inefficient. Current approaches (Tang et al., 2025; Zhang et al., 2025b) typically use the same LLM for reasoning and generation, forcing each output token generation to wait for all preceding reasoning tokens. Non-reasoning training can be accelerated by feeding all ground-truth tokens simultaneously, enabling parallel optimization. With reasoning tokens, this parallelism is lost, as the model must wait for the reasoning outputs, slowing down training. For inference, the time cost also increases with the total number of reasoning tokens. To solve the problem, the key is to keep the parallelism.
@@ 4.1 Experiment Settings
Datasets & Backbone LLMs. We evaluate our FlyThinker on three long-form personalized generation tasks from the user-based split of the LongLamp benchmark (Kumar et al., 2024): Product Review, Abstract Generation, and Topic Writing, which are designed to assess the effectiveness of LLMs in generating personalized, long-text outputs. The model is required to generate a long-form response that aligns with user preferences. Our primary backbone is Qwen2.5-3B-Instruct, with additional experiments on Qwen2.5-7B-Instruct and Gemma-7B-it available in Appendix L and  M.
Baselines & Evaluation Metrics. We compare FlyThinker against the following baselines: (1) Tuning-free methods: query-only generation (Non-pers(Yang et al., 2024), retrieval-augmented generation (RAG(Lei et al., 2023), and context steering (CoS(He et al., 2025); (2) Tuning-based methods: supervised fine-tuning (SFT(Salemi et al., 2024), summary-enhanced training (LLM-TRSR(Zheng et al., 2024), causal preference modeling (NextQuill(Zhao et al., 2025c), chain-of-thought reasoning (CoT(Wei et al., 2022; Salemi et al., 2025b), and latent reasoning (Coconut(Hao et al., 2024). We report ROUGE-1, ROUGE-L, BLEU, and METEOR scores. Baseline details and results on other evaluation metrics are provided in the Appendix D and E, respectively.
@@ 5 Related Work
Personalized Long-Form Generation. As the demand for complex content—such as customized reports, creative narratives, and personalized analyses—continues to grow, personalized long-form generation has emerged as a critical research frontier (Salemi et al., 2025b; Kumar et al., 2024; Salemi et al., 2025a; Salemi and Zamani, 2025). Unlike short-text generation, long-form tasks require LLMs to maintain consistency and personalization quality over extended sequences. This requirement is often challenged by the context drift problem, in which LLMs gradually lose alignment with user preferences as the text length increases (Kumar et al., 2024; Salemi et al., 2025b).
The recent benchmark LongLaMP (Kumar et al., 2024) has been proposed to systematically evaluate LLMs’ capabilities in personalized long-form generation. To address quality degradation, some approaches incorporate Retrieval-Augmented Generation (RAG), which injects personalized signals into the LLMs’ context (Mysore et al., 2023a; Kumar et al., 2024). REST-PG to improve the LLM’s ability to leverage personal data more effectively throughout the generation process (Salemi et al., 2025b).
Despite these advances, a critical gap remains: existing methods struggle to reason over implicit user preferences and dynamically adapt to evolving contexts in complex, real-world long-form scenarios. FlyThinker directly addresses this limitation by introducing a dynamic, token-level reasoning mechanism that continuously integrates personalized signals during generation. By enabling fine-grained, adaptive personalization rather than relying on static conditioning strategies, FlyThinker achieves deeper contextual alignment and more robust personalization across extended outputs.
Reasoning-Enhanced Generation. Reasoning-enhanced generation leverages the reasoning capabilities of LLMs to guide the production of high-quality responses. Widely adopted paradigms such as explicit reasoning (Wei et al., 2022) and latent reasoning (Hao et al., 2024) encourage LLMs to articulate intermediate reasoning steps. These reasoning paradigms have been adapted for personalization to capture implicit user preferences. REST-PG (Salemi et al., 2025b) employs a think-then-generate paradigm in which reasoning paths over personal context are generated via self-training to improve alignment. R2P (Luo et al., 2025a) utilizes a hierarchical reasoning template to guide large reasoning models in a think-then-generate framework for structured personalized outputs. RPM (Kim et al., 2025b) introduces personalization for black-box LLMs by constructing structured, user-specific reasoning paths and retrieving reasoning-aligned examples prior to producing the final response. TagPR (Jin et al., 2025) enhances personalization reasoning via semantically tagged reasoning chains and reinforcement learning over a personalization reward model. AlignXplore (Li et al., 2025a) infers personalized preferences from behavioral signals through extended inductive reasoning chains before downstream generation.
The aforementioned methods primarily rely on a static, one-shot reasoning process, in which the LLM performs reasoning once before generating the entire response. While effective for short-form tasks (Luo et al., 2025a; Kim et al., 2025b; Jin et al., 2025; Li et al., 2025a), this think-then-generate paradigm often struggles with long-form personalized generation because it fails to adapt to evolving contexts and dynamic preference shifts as the content unfolds. Furthermore, the sequential separation of reasoning and generation introduces significant latency, limiting real-time applicability (Salemi et al., 2025b).
In contrast, by employing a parallel architecture that synchronizes a reasoning model with a generation model, our proposed FlyThinker enables on-the-fly reasoning that evolves dynamically alongside the response. This approach not only mitigates the training and inference efficiency bottlenecks inherent in sequential paradigms but also significantly enhances the LLM’s capacity for fine-grained, dynamic reasoning over user preferences, resulting in superior personalization performance in complex long-form tasks.
@@ 6 Conclusion
Ethics Statement
This work complies with the ICLR Code of Ethics. No human or animal subjects were involved. All datasets (Product Review, Abstract Generation, Topic Writing) were used in accordance with their usage guidelines, without privacy violations or personally identifiable information. We are committed to maintaining transparency, fairness, and integrity in all aspects of this research.
Reproducibility Statement
We have made every effort to ensure the reproducibility of our results. The experimental setup is detailed in Section 4.1 and Appendix C. We also provide pseudocode for both training and inference (see Appendix B) to assist in reproducing our experiments. Moreover, the datasets in LongLamp are publicly available, ensuring consistent and reproducible evaluation. We believe these measures will enable other researchers to replicate our work and advance the field.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62525211) and the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Appendix C Implementation Details
In our study, we fine-tune the Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct model to validate our approach. All experiments, including FlyThinker and baseline models, are conducted on four NVIDIA A100 GPUs. The LLMs are loaded via HuggingFace Transformers (Wolf et al., 2019), and the training is implemented in Python 3.10 using the Verl training framework (Sheng et al., 2024).
@@ Appendix H Latent Reasoning Trajectory: Visualization and Analysis
To examine whether these trajectories relate to the preference, we used GPT-5.1 to compare the samples. We selected the two samples whose users had the most similar trajectory directions (sample0 and sample4) and the one farthest away (sample2) in Figure 7. GPT-5.1 judged sample4 to be most similar to sample0 in paragraph structure, narrative style, and analytical approach. Sample2 was the least similar. These results show that different users induce distinct reasoning trajectories, and that closer trajectories correspond to more similar writing styles and semantic structures, etc. This suggests that latent reasoning tokens may contain interpretable semantic information useful for user clustering and preference visualization.
@@ Appendix I Reasoner–Generator Relative Size Analysis in FlyThinker
In this section, we added an LLM-based evaluation to provide a more comprehensive assessment of personalized generation quality. Specifically, we use GPT-4o as an evaluator and compare the outputs produced by FlyThinker with those from the strongest baseline, SFT, following a structured evaluation rubric 9. As shown in the Table 5, FlyThinker consistently achieves higher scores across all datasets under the LLM-based evaluation. These findings indicate that FlyThinker’s personalization gains are reflected not only in traditional automatic metrics but also in the LLM judge’s assessments.
@@
Appendix Q Difficulty-Aware Evaluation for Personalized Tasks